OpenStudy (anonymous):

MATH challenge Part 2

OpenStudy (lxelle):

waiting.

OpenStudy (anonymous):

yes wait

OpenStudy (lxelle):

Is this a troll? Lol.

OpenStudy (anonymous):

Give thema challeng from Priniciples of QM by P.A.M DIrac and all will fall down to their knees

OpenStudy (anonymous):

\[\huge \frac{ \sqrt{12}-\sqrt{27} }{ 6 }*\sqrt{3}*\sqrt{4}\] EVALUATE WITHOUT A CALCULATOR IF YOU REALLY HAVE THE GUTS DO IT WITHOUT GOOGLE

OpenStudy (rational):

-1

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\sqrt{12} - \sqrt{27}/6 * \sqrt{12}\] \[\frac{ \sqrt{12}(\sqrt{12} - \sqrt{27} )}{ 6 } => \frac{ 12 - 18}{ 6} => \frac{ -6 }{ 6 } => \]

OpenStudy (anonymous):

YUP!

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so what's difficult in this so that you are calling it a 'challenge'?

OpenStudy (rational):

\[\large \frac{ \sqrt{12}-\sqrt{27} }{ \sqrt{3} }= \frac{\sqrt{3} \left( \sqrt{4}-\sqrt{9} \right) }{ \sqrt{3} } = 2-3 = -1\]

OpenStudy (anonymous):

lol sorry it was easy lol , I got a 324 in the root and didn't get a square of what it is so thought it is diifficult sorry

OpenStudy (anonymous):

Do you have other questions ??

OpenStudy (anonymous):

I am solving , i will post if i have intelligent ones ,

OpenStudy (anonymous):

this was a stupid question sorry every1

OpenStudy (anonymous):

what is your grade??

OpenStudy (anonymous):

now if upost such challenge please do come with some hard one...what about matrices they still proves hard

OpenStudy (anonymous):

and so long